Calculate how many grams of KO2LAiD are needed to make 43.0 ml of 0.25 M solution
1
Expert's answer
2020-05-19T08:47:07-0400

0.25M means 0.25moles per 1000ml.

Then (43/1000)* 0.25 = 0.01075 moles are in 43ml of solution."Mass = moles*rfm"

Mass =0.01075 *71.1 = 0.76g